One train done distance of 60km at 60km per hour. and return time same distance but speed is 40km per hoour. what is the average speed of the train?
Found 2 solutions by sachi, Cromlix:
Answer by sachi(548)

ongoing travel
the distanc= 60 km & speed =60 kmph
so the time=dist/speed=60/60=1 hr
return journey
the distanc= 60 km & speed =40 kmph
so the time=dist/speed=60/40=1.5 hrnow the total dist=60+60=120 km
& total time=1+1.5=2.5 hr
so the av speed =totaldist/total time=120/2.5=48 kmph

Answer by Cromlix(4381)

Total Average Speed = (2 * Speed 1 * Speed 2)/ (Speed 1 + Speed 2)
= ( 2* 60 * 40)/ (60 + 40)
= 4800/100
= 48 kmph
Hope this helps.
:-)